OpenStudy (anonymous):

Suppose the population of a town is 100,000 in 1999. The population increases at a rate of 4.5% each year. What will be the population of the town in 2005?

OpenStudy (anonymous):

If you had an equation of the rate of change, then you could solve with dy/dt, but you could do this by multiplying by .045 and add it to the original value I guess. Afraid I don't remember the exact formula

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

We're going to use the formula A = P(1+r)^t and this time the formula is correct because the population is increasing 5% each year.

jimthompson5910 (jim_thompson5910):

In this case, P = 100000, r = 0.045 and t = 6 (because 6 years have passed from 1999 to 2005)

jimthompson5910 (jim_thompson5910):

So we go from A = P(1+r)^t to A = 100000(1+0.045)^6

jimthompson5910 (jim_thompson5910):

what happens next?

OpenStudy (anonymous):

A= 130,226

jimthompson5910 (jim_thompson5910):

you nailed it

jimthompson5910 (jim_thompson5910):

very nice job

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome